3- . Bond X is an 8% semi-annual coupon bond with a par value of $1000 and a maturity of 10 years. The bond has a YTM of 7%. What is the value of the bond?

4. Bond J is a 10% semi-annual coupon bond with a par value of $1000 and a maturity of 2 years. If the assumed spot rates for a two year period are as follows, what is the value of the bond?

Maturity (in years)

Spot rates

0.5

3.79%

1.0

3.93%

1.5

4.12%

2.0

4.34%

5. If a zero coupon bond is currently trading at $75.46 and has a par value of $100 and a maturity of 4 years, what is its yield to maturity?
Answer #1

1.
=1000*8%/7%*(1-1/1.035^20)+1000/1.035^20=1071.06201650976

2.
=5%*1000/(1+3.79%/2)+5%*1000/(1+3.93%/2)^2+5%*1000/(1+4.12%/2)^3+5%*1000/(1+4.34%/2)^4+1000/(1+4.34%/2)^4=1107.79226217257

3.
=(100/75.46)^(1/4)-1=7.29285%